1. The inverse demand for leather is given by P = 50 - 0.5Q. The industry supply of leather is determined by its marginal cost: MC = 0.4Q. Unfortunately, the production of leather casues noxious chemical residue to leach into groundwater supplies. The external marginal cost caused by these residues grows with the amount of output, and is measured as EMC = 0.05Q. (Graph the inverse demand curve, private marginal cost, and social marginal cost curves.)

1a) How many leather is produced in the free market if the externality is not corrected?

1b) What is the free market price of leather if the externality is not corrected?

1c) what is the social marginal cost?

2. Now suppose that the government wishes to reduce the externality to efficient levels by levying a tax on leather production. (Graph the amount of tax, tax revenue collected by the government and DWL.)

2a) How much should that tax need to be?

2b) What is the resulting net price paid by buyers once the tax is in place?

2c) What is the resulting price received by the sellers once the tax is in place?

2d) Who paid more of the tax? Why? Explain.

Have an exam soon, therefore I need the proper steps for each please! Thank you!
Answer #1

a).

Consider the given problem here the inverse demand curve is given by, “P = 50 - 0.5*Q” with “MC = 0.4*Q”. So, at the equilibrium “P = MC”.

=> P = MC, => 50 – 0.5*Q = 0.4*Q, => 50 = 0.9*Q, => Q = 50/0.9 = 55.56. So, here the total leather production is given by, “Q=55.56”.

b).

Now, the price is given by, “MC = 0.4*Q”, => “MC = 0.4*55.56 = 22.22”. So, the free market price is given by, “P = 22.22”.

c).

Now, the social marginal cost is given by the sum of “private marginal cost” and the “external cost”, => SMC = PMC + EC.

=> SMC = MC + EC = 0.4*Q + 0.05*Q = 0.45*Q = SMC.

So, the socially optimum level of output will be determined by the intersection of “SMC” and the “P”. Consider the following fig.

SMC PMC+EC PMC E2 23.68 E1 22.23 52.63 55.56

At the equilibrium “P=SMC”.

=> 50 – 0.5*Q = 0.45*Q, => 50 = 0.95*Q, => Q = 50/0.95 = 52.63 and the corresponding price is given by “P = 50 – 0.5*Q = 50 – 0.5*52.63 = 23.68 = P.

So, here “E1” be the free market equilibrium and “E2” be the socially optimum equilibrium.